Find the value of r in the geometric sequence 200, 20, 2, 0.2
Answer by mananth(16946)

Find the value of r in the geometric sequence 200, 20, 2, 0.2
..
0.2/2=1/10
2/20=1/10
......
r=1/10
m.ananth@hotmail.ca